waht is the answer to this problem 2r-s=1 and 3r+2s=19

Answer by Nate(3500)

2r-s=1
and
3r+2s=19
First, get s by itself
2r-s=1
-s=-2r+1
s=2r-1
plug 2r-1 for s in the other equation
3r+2s=19
3r+2(2r-1)=19
3r+4r-2=19
7r=21
r=3
plug in r in one of the equations
2r-s=1
6-s=1
-s=-5
s=5